How to calculate ratio of specific heat in adiabatical process (TV^y).how?

Grade:11

2 Answers

Arun
5 years ago
You want to have ratio of Cp and Cv.
 
Then just use the relationship Cp h Cv = nR
and
Its ratio is so defined in terms of gamma which is guven
Aastha Singh
5 years ago
DEAR STUDENT,
ENTHALPY TO INCREASE TEMPERATURE BY ^T AT CONSTANT PRESSURE, DELTA H= nCP^T
ENTHALPY TO INCREASE TEMP. BY ^T AT CONSTANT VOLUME, DELTA U= nCv^T
RATIO OF SPECIFIC HEAT = CP/CV= GAMMA
APPLY PV=nRT , TO EQUATE TO ADIABATIC CONDITION